# Formation and disruption of wide binaries in star clusters revealed by N-body simulations

### 02 Apr 2026 v3

@@ Abstract
Wide (soft) binaries are expected to be rapidly disrupted in dense stellar environments, yet they are observed in both the Galactic field and open clusters (OCs). In this paper, we investigate the formation and disruption of wide binaries in star clusters using direct NN-body simulations. We perform simulations containing 10,000 objects with varying binary fractions and initial bulk rotation to give an in-depth look into the dynamical evolution of wide binaries in star clusters. We find that wide binaries dominate early disruption and formation processes during the initial high-density phase of cluster evolution. We propose two semi-analytical models to reproduce the evolution of the wide-binary population in simulations. The exponential model consists of an early, rapid-disruption phase with a time less than 10 Myr, driven by frequent encounters at high density, and a longer, relaxation-driven phase between 200 and 300 Myr. The broken power-law model provides break timescales when the decrease of wide binaries slows down during the early and long-term disruption. All timescales from both models agree with each other and decrease with increasing stellar density induced by high primordial binary fraction and cluster rotation. Wide binary disruption is mostly responsible for the early decline in the total binary fraction of the cluster. Such disruption leads to the decrease of radial binary fraction toward the cluster center until 500 Myr. Our results suggest low-density OCs or stellar groups younger than 10 Myr as the optimal environments for detecting wide binaries and provide a physical framework for understanding their contribution to the Galactic field population.
@@ I Introduction
Most stars are observed in multiple systems (Duchêne and Kraus, 2013; Offner et al., 2023). Among them, binary systems are the fundamental components of stellar multiplicity and play a crucial role in regulating the dynamical evolution of stellar systems, such as star clusters. In dense environments where close encounters are frequent, binaries are commonly classified according to their binding energy: hard binaries, which possess high binding energies and statistically tend to harden through interactions, and soft binaries, which have low binding energies and statistically are progressively softened (Heggie, 1975a; Hills, 1975). Soft/wide binaries have large semi-major axis and small binding energies, which leads to disruption before the relaxation timescale (Heggie, 1975b).
Despite their tendency to disrupt, wide binaries are still observed in the Galactic field and open clusters (OCs). El-Badry and Rix (2018) found more than 53,000 candidates with projected separation of 50–50,000 au within 200 pc via Gaia DR2. Similar study with Gaia eDR3 (El-Badry et al., 2021) identified 1.1 million binaries with 99% confidence, more than 50% of which have separation $\geq 1\,000$ au. Past studies suggested that wide binaries observed in the field are dynamically formed and have outlived their natal clusters (Kouwenhoven et al., 2010; Moeckel and Bate, 2010; Moeckel and Clarke, 2011; Perets and Kouwenhoven, 2012).
@@
Previous studies of star clusters have explored a wide range of formation and evolutionary processes. In systems containing binaries, numerous works have investigated how the binary fraction depends on cluster properties and membership (Ivanova et al., 2005; Hurley et al., 2007; Fregeau et al., 2009; Bonatto et al., 2012; Brinkmann et al., 2017). Other studies have focused on cluster rotation and the role of primordial binaries in shaping rotational evolution and angular momentum transport (Einsel and Spurzem, 1999; Spurzem, 1999; Hong et al., 2013; Mapelli, 2017; Bianchini et al., 2018; Livernois et al., 2022; Kamlah et al., 2022b; Tiongco et al., 2022; White et al., 2025). However, most of the aforementioned works included mainly hard binaries and excluded wide binaries. Some attempts on soft binary simulations are currently in progress (Wu et al., 2025; Yang et al., 2026).
In this paper, we are motivated to explore the formation and disruption of wide binaries in the clustered environment by NN-body simulations. The evolution of wide binaries in star clusters is investigated under various initial conditions, e.g., primordial binary fraction and initial bulk rotation. We aim to quantify the evolution of wide binaries in star clusters via an analytical model to estimate the disruption timescales.
In Section II, we introduce our NN-body simulations. In Section III, the results of our simulations of OCs with different primordial binary fractions and rotation rates are presented. We discuss binary disruption/formation in Section III.1, and propose two semi-analytical models for wide binary evolution (Section III.2). The evolution of binary fraction and its radial dependence are examined in Section III.3. In Section IV, we summarize our results.
II NN-body simulations
We use NBODY6++GPU, a high-performance GPU-accelerated code for gravitational NN-body simulations (Kamlah et al., 2022a; Spurzem and Kamlah, 2023). NBODY6++GPU is a successor of the direct force integration NN-body codes originally written by Sverre Aarseth (Aarseth, 1985, 1999, 2003; Spurzem, 1999; Aarseth et al., 2008).
Descriptions of the models are shown in Table 1. There are 13 sets of NN-body simulations with different parameters. We use McLuster (Küpper et al., 2011; Kamlah et al., 2022a; Leveque et al., 2022) to generate initial models, and the resulting outputs are used to run all simulations. The density model is based on the King model (King, 1966), with a concentration parameter ($W_{0}$) of 6.0. To ensure robustness of the results, we run 5 to 10 models ($n$ in Table 1) with the same initial conditions but different random seeds to reduce stochasticity.
@@ II NN-body simulations
All simulations have the same number of objects ($N_{obj}$) of 10,000, including both singles and binaries. A binary system is considered a single object. Therefore, the number of stars, NN, differs from the number of objects. The binary fraction ($f_{b}$) is defined as the ratio of the number of binary systems ($N_{b}$) over $N_{obj}$ ($f_{b}=N_{b}/N_{obj}$). We initialize hard and soft primordial binaries with differing setups. For hard binaries, the semi-major axis ($a$) is set to follow a uniform distribution in $\rm log(a)$ in the range of 0500-50 au for $M<5M_{\odot}$ and Sana et al. (2012) distribution for $M>5M_{\odot}$. For soft binaries, we use the right triangular distribution (also known as ‘flat uniform distribution’ in McLuster): $f_{x}=2(a-a_{min})/(a_{max}-a_{min})$, where $a_{min}=50$ au and $a_{max}=50,000$ au. The choice of this distribution for soft binaries is to maximize the number of wide binaries. The ratio of hard over soft binaries is $1/4$. We also divide the binaries in our models into 3 distinct groups by their semi-major axis separation: close ($a\leq 10$au), intermediate ($10<a\leq 1000$au), and wide binaries (a>1000a>1000au).
Table 1: Models parameters and fitted timescales
Category Name $n$ N $f_{\rm b,h}$ [%] $f_{\rm b,s}$[%] $f_{\rm b,t}$ [%] $\omega_{0}$ $t_{1}$ [Myr] $t_{2}$ [Myr] $t_{\rm br1}$ [Myr] $t_{\rm br2}$ [Myr]
@@
Note. — $n$ denotes the number of models in each set of simulation. NN is the initial number of all the stars in each model. $f_{\rm b,h}$ and $f_{\rm b,s}$ are the binary fraction of hard and soft binaries. $f_{\rm b,t}$ is the total binary fraction of hard and soft binaries. $\omega_{0}$ is the dimensionless angular velocity. $t_{1}$ and $t_{2}$ are the short and long-term wide binary dissolution timescales of equation (1). $t_{\rm br1}$ and $t_{\rm br2}$ are break timescales when the decrease of wide binaries slows down from equation (2)
We divide our models into several categories, such as Binary influence (BI), rotation, and BI on rotation. In BI models, we investigate the cluster evolution with different $f_{b}$. For rotating models, we use FOPAX (Einsel and Spurzem, 1999; Kim et al., 2002, 2004, 2008) to generate the 2D orbit-averaged Fokker-Planck initial models that add rotation to NN-body simulations. Further, we combine them with the McLuster output with proper NN-body format scaling via the Monte Carlo rejection technique as described in previous similar studies (Kamlah et al., 2022b; Bissekenov et al., 2025). Rotating King models are parametrized by pair $(W_{0},\omega_{0})$, where $\omega_{0}$ is dimentionless angular velocity (Table 1, Einsel and Spurzem, 1999; Kamlah et al., 2022b; Bissekenov et al., 2025). Higher value of $\omega_{0}$ indicates higher level of rotation.
@@ III.1 Binary disruptions/formations
We first look into the evolution of single stars and binaries in the modeled clusters. We use the set of simulations N15k_10hb_40sbN15k\_10hb\_40sb (Table 1) as an example to display the results throughout this paper. Figure 1 shows the evolution of the total number of stars, single stars, and binaries (panel (a)) and the number of close, intermediate, and wide binaries (panel (b)). The simulation is initialized with $f_{bin}=50\%$, meaning 5,000 binaries, but in panel (a), the initial number of binaries is $\approx 2000$. This is due to the instantaneous disruption of very wide binaries caused by the calculations of binding energy in the code. Pairs in disrupted binaries are counted as single stars.
@@
To gain a more in-depth understanding of binary disruptions and formations, we record dynamical events in binaries, labeled as disruption, formation, and escaper. Figure 2 shows the evolution of the number (panel (a)), semi-major axis distribution (panel (b)), and eccentricity distribution (panel (c)) for the binaries that were disrupted, formed, and escaped during the simulation. Wide binaries have binding energies very close to zero, making them appear temporarily disrupted (binding energy >0>0) in the internal regularizations of NN-body codes because of stellar flybys or passages through the cluster center. To ensure physical consistency, we recompute the binding energies for all recorded events and keep only those disruptions with positive binding energy, together with their corresponding formation events. All other temporary (repeated) disruption/formation events are removed from the analysis of the data in panels (a), (b), and (c) in Figure 2, except panel (d) where we show one specific binary experiencing those temporary events. We also do not include instantaneous disruptions at age 0 Myr.
As shown in panel (a) of Figure 2, most disruption and formation events occur at early stages of cluster evolution. These are the disruptions of primordial soft binaries, which are vulnerable to perturbation due to their wide separation (panel (b) in Figure 2). As can be seen in panel (b), wide binaries are significantly dominant in the disruption and formation events ($a>1000$ au), compared to intermediate and close binaries ($a<1000$ au). Their evolution is easily affected by encounters, and therefore, they are frequently disrupted. At the same time, formation follows a trend similar to disruption, but much less efficiently. This trend changes at later evolutionary stages, when the number of disruptions becomes comparable to the number of dynamical formations. Some of these late-time formations may be driven by three-body encounters (Goodman and Hut, 1993; Ginat and Perets, 2024; Atallah et al., 2024). The aforementioned effects collectively explain the overall decline of wide/soft binaries in Figure 1. Finally, all event types exhibit a thermal eccentricity distribution in panel (c) of Figure 2, as expected from the initially imposed thermal distribution.
As for escapers, they are mostly close binaries and they start escaping from the cluster after 20 Myr. On the contrary, wide binaries have less chance to escape since the disruption timescale is much shorter than the escaping timescale (panel (b)). These behaviors of binaries are consistent with the studies of Heggie (1975a, b) and Hills (1975).
@@
Figure 2: Events of disruption (red), formation (blue), and escape (orange) for all binaries as a function of time (a), distributions of semi-major axis (b), and eccentricity (c) for the set N15k_10hb_40sbN15k\_10hb\_40sb from Table 1. The age axis in (a) is logarithmic, early-time bins span much shorter intervals than those at later ages. (d) shows the evolution of actual separation (orange) and semi-major axis (blue) as a function of time of specific case of very wide binary with chaotic evolution or highly perturbed by stellar encounters. Temporary disruption and formation events during the evolution of wide binaries of this type were excluded from (a), (b), and (c).
@@ III.2 Wide binary evolution model
By analyzing the behavior of wide binaries presented in Section  III.1, we consider two semi-analytical models for the evolution of wide binaries in NN-body simulations.
@@
We first consider a two-part exponential model
@@
Another model we propose is a broken power law
@@
where $\alpha_{1}$ and $\alpha_{2}$ are the decrease rate of wide binaries and $t_{\rm br1}$ and $t_{\rm br2}$ are break timescales when the decrease of wide binaries slows down. We find that when we set $\alpha_{1}$ and $\alpha_{2}$ as variables, they converge to $\alpha_{1}=1.5$ and $\alpha_{2}=2$. Thus, we set them as constants.
@@
Performance of the models was quantified using three standard metrics: the coefficient of determination ($R^{2}$, Montgomery and Runger (2019)), which measures the proportion of variance in the data explained by the model; the root-mean-squared error (RMSERMSE, Chai and Draxler (2014)), which reflects the typical magnitude of prediction errors with greater sensitivity to large deviations; and the mean absolute error (MAEMAE, Chai and Draxler (2014)), which indicates the average absolute discrepancy between predictions and observations and is robust to outliers. The exponential and power-law models yield $R^{2}\approx 0.970$ and $0.988$, respectively (values closer to 1 indicate a better fit). Their RMSERMSE values are $\approx 7.98$ and $\approx 4.96$, and their MAEMAE values are $\approx 4.06$ and $\approx 1.99$. Both RMSERMSE and MAEMAE are measured in the number of wide binaries. Given the initial 4000 wide binaries, the values of RMSERMSE and MAEMAE are less than 1% and therefore acceptable.
@@
The fitted values of $t_{1}$ and $t_{2}$ of exponential model as well as $t_{\rm br1}$ and $t_{\rm br2}$ of the power-law model agree with each other respectively. We present their values for all the simulations in Table  1. Wide binary disruptions begin at the crossing time $t_{\rm cross}$, followed by a steady decline until $t_{1}$ ($t_{\rm br1}$). After that, the disruption rate slows and is driven mainly by encounters from two-body relaxation, decreasing further by $t_{2}$ ($t_{\rm br2}$). A final break occurs around $t_{2}$ ($t_{\rm br2}$), when only $\approx 5\%$ of the initially surviving wide binaries remain. The two models complement each other in explaining the wide binary disruption.

### 14 Apr 2026 v4

@@ Abstract
Wide (soft) binaries are expected to be rapidly disrupted in dense stellar environments, yet they are observed in both the Galactic field and open clusters (OCs). In this Letter, we investigate the formation and disruption of wide binaries in star clusters using direct $N$-body simulations. We perform simulations containing 10,000 objects with varying binary fractions and initial bulk rotation to give an in-depth look into the dynamical evolution of wide binaries in star clusters. We find that wide binaries dominate early disruption and formation processes during the initial high-density phase of cluster evolution. We propose two semianalytical models to reproduce the evolution of the wide-binary population in simulations. The exponential model consists of an early, rapid-disruption phase with a time less than 10 Myr, driven by frequent encounters at high density, and a longer, relaxation-driven phase between 200 and 300 Myr. The broken power-law model provides break timescales when the decrease of wide binaries slows down during the early and long-term disruption. All timescales from both models agree with each other and decrease with increasing stellar density induced by high primordial binary fraction and cluster rotation. Wide binary disruption is mostly responsible for the early decline in the total binary fraction of the cluster. Such disruption leads to the decrease of radial binary fraction toward the cluster center until 500 Myr. Our results suggest low-density OCs or stellar groups younger than 10 Myr as the optimal environments for detecting wide binaries and provide a physical framework for understanding their contribution to the Galactic field population.
@@ I Introduction
Most stars are observed in multiple systems (Duchêne and Kraus, 2013; Offner et al., 2023). Among them, binary systems are the fundamental components of stellar multiplicity and play a crucial role in regulating the dynamical evolution of stellar systems, such as star clusters. In dense environments where close encounters are frequent, binaries are commonly classified according to their binding energy: hard binaries, which possess high binding energies and statistically tend to harden through interactions, and soft binaries, which have low binding energies and statistically are progressively softened (Heggie, 1975a; Hills, 1975). Soft/wide binaries have large semi-major axes and small binding energies, which leads to disruption before the relaxation timescale (Heggie, 1975b).
Despite their tendency to disrupt, wide binaries are still observed in the Galactic field and open clusters (OCs). El-Badry and Rix (2018) found more than 53,000 candidates with projected separation of 50–50,000 au within 200 pc via Gaia DR2. A similar study with Gaia eDR3 (El-Badry et al., 2021) identified 1.1 million binaries with 99% confidence, more than 50% of which have separation $\geq 1000$ au. Past studies suggested that wide binaries observed in the field are dynamically formed and have outlived their natal clusters (Kouwenhoven et al., 2010; Moeckel and Bate, 2010; Moeckel and Clarke, 2011; Perets and Kouwenhoven, 2012).
@@
Previous studies of star clusters have explored a wide range of formation and evolutionary processes. In systems containing binaries, numerous works have investigated how the binary fraction depends on cluster properties and membership (Ivanova et al., 2005; Hurley et al., 2007; Fregeau et al., 2009; Bonatto et al., 2012; Brinkmann et al., 2017). Other studies have focused on cluster rotation and the role of primordial binaries in shaping rotational evolution and angular momentum transport (Einsel and Spurzem, 1999; Spurzem, 1999; Hong et al., 2013; Mapelli, 2017; Bianchini et al., 2018; Kamlah et al., 2022b; Livernois et al., 2022; Tiongco et al., 2022; White et al., 2025). However, most of the aforementioned works included mainly hard binaries and excluded wide binaries. Some attempts on soft binary simulations are currently in progress (Wu et al., 2025; Yang et al., 2026).
In this Letter, we are motivated to explore the formation and disruption of wide binaries in the clustered environment by $N$-body simulations. The evolution of wide binaries in star clusters is investigated under various initial conditions, e.g., primordial binary fraction and initial bulk rotation. We aim to quantify the evolution of wide binaries in star clusters via an analytical model to estimate the disruption timescales.
In Section II, we introduce our $N$-body simulations. In Section III, the results of our simulations of OCs with different primordial binary fractions and rotation rates are presented. We discuss binary disruption/formation in Section III.1, and propose two semianalytical models for wide binary evolution (Section III.2). The evolution of binary fraction and its radial dependence are examined in Section III.3. In Section IV, we summarize our results.
II $N$-body simulations
We use NBODY6++GPU, a high-performance GPU-accelerated code for gravitational $N$-body simulations (Kamlah et al., 2022a; Spurzem and Kamlah, 2023). NBODY6++GPU is a successor of the direct force integration $N$-body codes originally written by Sverre Aarseth (Aarseth, 1985, 1999, 2003; Spurzem, 1999; Aarseth et al., 2008).
Descriptions of the models are shown in Table 1. There are 13 sets of $N$-body simulations with different parameters. We use McLuster (Küpper et al., 2011; Kamlah et al., 2022a; Leveque et al., 2022) to generate initial models, and the resulting outputs are used to run all simulations. The density model is based on the King model (King, 1966), with a concentration parameter ($W_{0}$) of 6.0. To ensure robustness of the results, we run 5 to 10 models ($n$ in Table 1) with the same initial conditions but different random seeds to reduce stochasticity.
@@ II $N$-body simulations
All simulations have the same number of objects ($N_{\rm obj}$) of 10,000, including both singles and binaries. A binary system is considered a single object. Therefore, the number of stars, $N$, differs from the number of objects. The binary fraction ($f_{b}$) is defined as the ratio of the number of binary systems ($N_{b}$) over $N_{\rm obj}$ ($f_{b}=N_{b}/N_{\rm obj}$). We initialize hard and soft primordial binaries with differing setups. For hard binaries, the semi-major axis ($a$) is set to follow a uniform distribution in $\rm log(a)$ in the range of $\rm 0-50$ au for $M<5M_{\odot}$ and Sana et al. (2012) distribution for $M>5M_{\odot}$. For soft binaries, we use the right triangular distribution (also known as ‘flat uniform distribution’ in McLuster): $f_{x}=2(a-a_{\rm min})/(a_{\rm max}-a_{\rm min})$, where $a_{\rm min}=50$ au and $a_{\rm max}=50,000$ au. The choice of this distribution for soft binaries is to maximize the number of wide binaries. The ratio of hard over soft binaries is $1/4$. We also divide the binaries in our models into three distinct groups by their semi-major axis separation: close ($a\leq 10$au), intermediate ($10<a\leq 1000$au), and wide binaries (a>1000a>1000au).
Table 1: Model parameters and fitted timescales
Category Name $n$ N $f_{\rm b,h}$ $f_{\rm b,s}$ $f_{\rm b,t}$ $\omega_{0}$ $t_{1}$ $t_{2}$ $t_{\rm br1}$ $t_{\rm br2}$
(%) (%) (%) (Myr) (Myr) (Myr) (Myr)
@@
Note. — $n$ denotes the number of models in each set of simulation. $N$ is the initial number of all the stars in each model. $f_{\rm b,h}$ and $f_{\rm b,s}$ are the binary fractions of hard and soft binaries. $f_{\rm b,t}$ is the total binary fraction of hard and soft binaries. $\omega_{0}$ is the dimensionless angular velocity. $t_{1}$ and $t_{2}$ are the short and long-term wide binary dissolution timescales of equation (1). $t_{\rm br1}$ and $t_{\rm br2}$ are break timescales when the decrease of wide binaries slows down from equation (2)
We divide our models into several categories, such as binary influence (BI), rotation, and BI on rotation. In BI models, we investigate the cluster evolution with different $f_{b}$. For rotating models, we use FOPAX (Einsel and Spurzem, 1999; Kim et al., 2002, 2004, 2008) to generate the 2D orbit-averaged Fokker-Planck initial models that add rotation to $N$-body simulations. Further, we combine them with the McLuster output with proper $N$-body format scaling via the Monte Carlo rejection technique as described in previous similar studies (Kamlah et al., 2022b; Bissekenov et al., 2025). Rotating King models are parameterized by pair $(W_{0},\omega_{0})$, where $\omega_{0}$ is dimensionless angular velocity (Table 1, Einsel and Spurzem, 1999; Kamlah et al., 2022b; Bissekenov et al., 2025). A higher value of $\omega_{0}$ indicates higher level of rotation.
@@ III.1 Binary Disruptions/Formations
We first look into the evolution of single stars and binaries in the modeled clusters. We use the set of simulations N15k_10hb_40sbN15k\_10hb\_40sb (Table 1) as an example to display the results throughout this paper. Figure 1 shows the evolution of the total number of stars, single stars, and binaries (panel (a)) and the number of close, intermediate, and wide binaries (panel (b)). The simulation is initialized with $f_{b}=50\%$, meaning 5000 binaries, but in panel (a), the initial number of binaries is $\approx 2000$. This is due to the instantaneous disruption of very wide binaries caused by the calculations of binding energy in the code. Pairs in disrupted binaries are counted as single stars.
@@
To gain a more in-depth understanding of binary disruptions and formations, we record dynamical events in binaries, labeled as disruption, formation, and escaper. Figure 2 shows the evolution of the number (panel (a)), semi-major axis distribution (panel (b)), and eccentricity distribution (panel (c)) for the binaries that were disrupted, formed, and escaped during the simulation. Wide binaries have binding energies very close to zero, making them appear temporarily disrupted (binding energy >0>0) in the internal regularizations of $N$-body codes because of stellar flybys or passages through the cluster center. To ensure physical consistency, we recompute the binding energies for all recorded events and keep only those disruptions with positive binding energy, together with their corresponding formation events. All other temporary (repeated) disruption/formation events are removed from the analysis of the data in panels (a), (b), and (c) in Figure 2, except panel (d), where we show one specific binary experiencing those temporary events. We also do not include instantaneous disruptions at age 0 Myr.
As shown in panel (a) of Figure 2, most disruption and formation events occur at early stages of cluster evolution. These are the disruptions of primordial soft binaries, which are vulnerable to perturbation due to their wide separation (panel (b) in Figure 2). As can be seen in panel (b), wide binaries are significantly dominant in the disruption and formation events ($a>1000$ au), compared to intermediate and close binaries ($a<1000$ au). Their evolution is easily affected by encounters, and therefore, they are frequently disrupted. At the same time, formation follows a trend similar to disruption, but much less efficiently. This trend changes at later evolutionary stages, when the number of disruptions becomes comparable to the number of dynamical formations. Some of these late-time formations may be driven by three-body encounters (Goodman and Hut, 1993; Atallah et al., 2024; Ginat and Perets, 2024). The aforementioned effects collectively explain the overall decline of wide/soft binaries in Figure 1. Finally, all event types exhibit a thermal eccentricity distribution in panel (c) of Figure 2, as expected from the initially imposed thermal distribution.
As for escapers, they are mostly close binaries, and they start escaping from the cluster after 20 Myr. On the contrary, wide binaries have less chance to escape since the disruption timescale is much shorter than the escaping timescale (panel (b)). These behaviors of binaries are consistent with the studies of Heggie (1975a, b) and Hills (1975).
@@
Figure 2: Events of disruption (red), formation (blue), and escape (orange) for all binaries as a function of time (a), distributions of semi-major axis (b), and eccentricity (c) for the set N15k_10hb_40sbN15k\_10hb\_40sb from Table 1. The age axis in (a) is logarithmic; early-time bins span much shorter intervals than those at later ages. Panel (d) shows the evolution of actual separation (orange) and semi-major axis (blue) as a function of time for a specific case of very wide binary with chaotic evolution or highly perturbed by stellar encounters. Temporary disruption and formation events during the evolution of wide binaries of this type were excluded from (a), (b), and (c).
@@ III.2 Wide binary evolution model
By analyzing the behavior of wide binaries presented in Section  III.1, we consider two semianalytical models for the evolution of wide binaries in $N$-body simulations.
@@
We first consider a two-part exponential model:
@@
Another model we propose is a broken power law:
@@
where $\alpha_{1}$ and $\alpha_{2}$ are the decrease rate of wide binaries, and $t_{\rm br1}$ and $t_{\rm br2}$ are break timescales when the decrease of wide binaries slows down. We find that when we set $\alpha_{1}$ and $\alpha_{2}$ as variables, they converge to $\alpha_{1}=1.5$ and $\alpha_{2}=2$. Thus, we set them as constants.
@@
Performance of the models was quantified using three standard metrics: the coefficient of determination ($R^{2}$, Montgomery and Runger (2019)), which measures the proportion of variance in the data explained by the model; the root-mean-squared error ($\rm RMSE$, Chai and Draxler (2014)), which reflects the typical magnitude of prediction errors with greater sensitivity to large deviations; and the mean absolute error ($\rm MAE$, Chai and Draxler (2014)), which indicates the average absolute discrepancy between predictions and observations and is robust to outliers. The exponential and power-law models yield $R^{2}\approx 0.970$ and $0.988$, respectively (values closer to 1 indicate a better fit). Their $\rm RMSE$ values are $\approx 7.98$ and $\approx 4.96$, and their $\rm MAE$ values are $\approx 4.06$ and $\approx 1.99$. Both $\rm RMSE$ and $\rm MAE$ are measured in the number of wide binaries. Given the initial 4000 wide binaries, the values of $\rm RMSE$ and $\rm MAE$ are less than 1% and therefore acceptable.
@@
The fitted values of $t_{1}$ and $t_{2}$ of the exponential model, as well as $t_{\rm br1}$ and $t_{\rm br2}$ of the power-law model, agree with each other, respectively. We present their values for all the simulations in Table 1. Wide binary disruptions begin at the crossing time $t_{\rm cross}$, followed by a steady decline until $t_{1}$ ($t_{\rm br1}$). After that, the disruption rate slows and is driven mainly by encounters from two-body relaxation, decreasing further by $t_{2}$ ($t_{\rm br2}$). A final break occurs around $t_{2}$ ($t_{\rm br2}$), when only $\approx 5\%$ of the initially surviving wide binaries remain. The two models complement each other in explaining the wide binary disruption.